OpenStudy (anonymous):

how do u find the x and y intercept from ... y=3x-6

OpenStudy (anonymous):

when solving for x, make y =0, and when solving for y, make x =0 X: 0=3x-6 6=3x 2=x (2,0) Y: y=3(0)-6 y=-6 (0,-6)

OpenStudy (anonymous):

wait wouldnt it be (2,-6)

OpenStudy (anonymous):

if you are finding the intercepts, then the y intercept is where the line crosses the axis, and therefore x must be equal to 0, and the opposite is also true.. therefore, you will have 2 points, (2,0) <--where the line crosses x axis, and (0,-6) <---where line crosses y axis

OpenStudy (anonymous):